OpenStudy (i<3mybaby):

In the triangle below, what is cos θ? A. twelve-thirteenths B. five-thirteenths C. thirteen-twelfths D. twelve-fifths

OpenStudy (i<3mybaby):

@jcpd910 @jdoe0001 @johnweldon1993 @iambatman @adrynicoleb @anonymous_user @jim_thompson5910 @mathmate

OpenStudy (adrynicoleb):

Sorry, I don't know :/

OpenStudy (i<3mybaby):

:(

OpenStudy (e.mccormick):

Know your trig ratios? SOH CAH TOA

OpenStudy (i<3mybaby):

Yes but I am just lost in steps and im like uggh

OpenStudy (e.mccormick):

Well, it is cosine, so CAH. CAH means C=A/H or: \(\cos\theta=\dfrac{\text{Adjacent}}{\text{Hypotenuse}}\)

OpenStudy (i<3mybaby):

ok

OpenStudy (e.mccormick):

So put the measure of the adj side over they hyp and then reduce the fraction to lowest terms.

OpenStudy (i<3mybaby):

its 5/13

OpenStudy (e.mccormick):

Opp and Adj mean Across and Next to. They Hypotenuse is always the one related to the right angle, so it never changes.  Make sure you have your Adj correct for that angle.

OpenStudy (i<3mybaby):

OK.... `